# Functional inequalities on the biased and restricted cube: an induction approach

### v2

@@ Abstract
We develop new discrete functional inequalities on the hypercube via the induction-by-restrictions method. This method reduces high-dimensional inequalities to explicit low-dimensional analytic verifications and has recently proved effective for many discrete functional inequalities. We establish two results along this framework. First, we prove a sharp $p$-biased edge-isoperimetric inequality for real-valued increasing functions, which recovers the classic biased edge-isoperimetric inequality for increasing sets and identifies increasing subcubes as the extremizers. This result also admits a probabilistic interpretation in terms of maximizing the mean first exit time of biased random walks. Second, we give an inductive proof of a Poincaré inequality on increasing subsets of the cube that was recently established by Fei and Ferreira Pinto Jr [17], yielding an $O(n^{2})$ upper bound on the mixing time of censored random walks, improving upon previous bounds.
@@ 1 Introduction
Functional inequalities on the discrete cube, such as the Poincaré, log-Sobolev, and edge-isoperimetric inequalities, form a central part of modern discrete analysis. They connect Boolean function analysis, discrete isoperimetric problems, and the spectral theory of Markov chains, providing powerful tools for studying concentration of measure, threshold phenomena, and mixing times of random walks. In the classical uniform product setting on $\{0,1\}^{n}$, these inequalities are well understood: sharp constants are known (e.g., the exact log-Sobolev constant [6]) and elegant proofs arise via tensorization [20], semigroup methods [23, 24], or discrete Fourier analysis [13]. However, once one departs from the product setting–by restricting to structured subsets or by working under biased measures–classical approaches often fail or lose sharpness. Recent progress highlights both the challenges and the opportunities in these non-product environments: FKN-type structure theorems and sharp junta thresholds for low-degree Boolean functions on the slice/Johnson graph [18], (near) sharp log-Sobolev inequalities on the multislice [19], and representation-theoretic analyses of low-degree functions on the symmetric group that yield asymptotically sharp edge-isoperimetric bounds for the transposition graph [14]. These developments motivate the search for new discrete functional inequalities adapted to non-product environments.
A case in point that illustrates both the promise and the limitations of classical tools is the following. On the uniform cube, the logarithmic Sobolev inequality is tight for general real-valued functions, yet when specialized to indicator functions it recovers the edge-isoperimetric bound only up to a suboptimal multiplicative constant (by a factor $1/\ln 2$). Samorodnitsky [30] overcame this mismatch by proving a discrete functional inequality tailored to subsets: for functions supported on $A\subseteq\{0,1\}^{n}$ it yields the sharp Harper’s constant and characterizes equality via codimension-$k$ subcubes. Although Samorodnitsky’s inequality is formulated on the hypercube, recovering Harper’s optimal constant forces the right–hand side to involve the conditional average over the support set $A$. This quantity depends on the global geometry/size of $A$ and is not additive over coordinates; it cannot be written as an average of one-coordinate functionals. Consequently, the standard tensorization machinery for product spaces, which relies on coordinatewise factorization, cannot “lock onto” the correct constant. In this sense, even though the ambient space is product, conditioning on an arbitrary $A$ unfortunately destroys tensor structure at the level of the inequality, revealing difficulties that mirror those in genuinely non-product settings. In the $p$-biased setting, while a sharp edge-isoperimetric inequality for increasing sets is known (see, e.g., [15, 25]), the natural $p$-biased log-Sobolev inequalities (e.g., [7, 9]) again fail to deliver the optimal constant for indicators except in limiting regimes (such as $p\to 0$), leaving a quantitative and conceptual gap. These considerations highlight an urgent need to establish a Samorodnitsky-type functional inequality on the $p$-biased cube.
A second illustration comes from mixing of censored random walks, where product structure is explicitly broken. If we censor the simple random walk on $\{0,1\}^{n}$ to a subset $A$ (only moves that stay in $A$ are accepted), the chain now lives on a non-product space with geometry dictated by the boundary of $A$. In this setting, standard product-based tools (tensorization, Fourier decompositions) no longer apply cleanly, and the behavior can vary widely: the walk may mix well, but it can also mix very poorly even when $A$ is large and connected (see [17, Examples 1–2]). Motivated by the question of Ding and Mossel [10, Question 1.1]—whether $t_{\mathrm{mix}}(\varepsilon)\leq O_{c,\varepsilon}(n\log n)$ holds for increasing sets of constant density—recent progress of Fei and Ferreira Pinto Jr [17] proves an $O(n^{2})$ bound by establishing a Poincaré inequality on the restricted (non-product) space. Their argument relies on a rather complex directed heat-flow method. This naturally raises a question: can simpler methods–such as an induction argument on the dimension–still be effective in such non-product spaces? Pursuing this direction can clarify which obstacles are intrinsic to the loss of product structure and offer a cleaner pathway toward functional inequalities in genuinely non-product settings.
A particularly powerful paradigm that has recently emerged for proving such inequalities is the induction by restrictions combined with two-point inequalities, an approach widely adopted within the combinatorics and discrete analysis community, as exemplified by [1, 2, 3, 8, 12, 15, 16, 25, 30, 33].
Our contributions. In this paper we establish two new discrete functional inequalities along this framework: one concerns a sharp $p$-biased edge-isoperimetric inequality for all real-valued increasing functions, and the other is a Poincaré inequality on increasing subsets of the hypercube. Both are proved by induction on the dimension combined with a two-point inequality, illustrating the versatility of this method across departures from the product setting.
  • (1)
    $p$-biased edge-isoperimetric inequality. In the uniform cube, Harper’s edge-isoperimetric theorem identifies subcubes as the unique minimizers of edge boundary, and Samorodnitsky [30] established a functional extension with sharp constants. For the $p$-biased hypercube endowed with product measure $\mu_{p}$, however, only weaker forms were known, typically losing sharp constants. We establish a new $p$-biased edge-isoperimetric inequality for real-valued increasing functions via induction on dimension and a two-point verification. Our result recovers the sharp $p$-biased isoperimetric inequality for increasing sets, and admits a probabilistic interpretation: increasing subcubes maximize the mean first exit time of the $p$-biased random walk among increasing sets of fixed measure.
  • (2)
    Poincaré inequality on increasing sets. Ding and Mossel [10] introduced the censored random walk on an increasing subset $A\subseteq\{0,1\}^{n}$ and asked whether it mixes in $O(n\log n)$ steps under constant density. They proved an $O(n^{3})$ upper bound. Fei and Ferreira Pinto Jr [17] established a Poincaré inequality on increasing sets that improves the spectral gap and implies an $O(n^{2})$ mixing-time upper bound. We give a simpler inductive proof of this Poincaré inequality, again leading to the $O(n^{2})$ mixing bound for the censored random walk.
1.1 Functional version of the biased edge-isoperimetric inequality
In this subsection we work on the $n$-dimensional hypercube. The state space is the graph $Q^{n}=(\{0,1\}^{n},E)$ with $2^{n}$ vertices indexed by binary strings, where two vertices are adjacent (written $x\sim y$) iff they differ in exactly one coordinate. Let $g:\{0,1\}^{n}\to\mathbb{R}$, we define the Dirichlet form of $g$ as
where the expectation is taken with respect to the uniform probability measure $\mu:=\mu_{1/2}^{\otimes n}$ on the hypercube. Furthermore, we denote by $x\sim A$ that $x$ is drawn uniformly at random from the subset $A\subseteq\{0,1\}^{n}$. In 2017, Samorodnitsky [30] established the following discrete functional inequality on the hypercube.
Theorem 1.1 (Samorodnitsky [30]).
A subcube of co-dimension $k$ is a set of the form
$\{x\in\{0,1\}^{n}:x_{i_{1}}=a_{1},x_{i_{2}}=a_{2},\dots,x_{i_{k}}=a_{k}\},$
where $i_{1}<\cdots<i_{k}$ are coordinate indices, and $a_{1},\dots,a_{k}$ are fixed elements of $\{0,1\}$. The coordinates indexed by $i_{1},\dots,i_{k}$ are called the fixed coordinates; the other coordinate indices are called the active coordinates.
Remark.
  • (1)
    When $g$ is an indicator function of $A$, the Samorodnitsky’s inequality (1) implies the classic edge-isoperimetric inequality on $Q^{n}$: for every $A\subset\{0,1\}^{n}$,
    $|\nabla A|\geq|A|\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right),$ (2)
    where the edge boundary $\nabla(A)$ is the set of edges between $A$ and its complement $A^{c}=V(Q^{n})\setminus A$. Equality in (2) holds if and only if $A$ is a subcube of co-dimension $k$, for any $0\leq k\leq n$. It is easy to see that $\mathcal{E}(\mathbbm{1}_{A},\mathbbm{1}_{A})=\frac{2|\nabla A|}{2^{n}}$. The fully edge isoperimetric problem for $Q^{n}$ was solved by Harper [21], Lindsey [28], Bernstein [5], and Hart [22], which states that among all subsets of $\{0,1\}^{n}$ of size $1\leq k\leq 2^{n}$, the first $k$ elements of the lexicographic ordering on $\mathcal{P}([n])$ (defined by $x<y\Leftrightarrow\max(x\Delta y)\in y$) have the smallest possible edge-boundary.
  • (2)
    A primary motivation for Samorodnitsky’s functional inequality is to obtain a general discrete functional inequality for real-valued functions that recovers the sharp/correct constant in (2). In contrast, the classical logarithmic Sobolev inequality on the hypercube only recovers (2) up to a multiplicative factor $\frac{1}{\ln 2}$:
    $\mathcal{E}(g,g)\geq 2\cdot{\rm Ent}(g^{2}):=2\mathbb{E}[g^{2}\ln(g^{2})]-2\mathbb{E}[g^{2}]\ln(\mathbb{E}[g^{2}]).$ (3)
  • (3)
    To see the connection between Samorodnitsky’s inequality (1) and the log-Sobolev inequality (3), we rewrite it as
    $\mathcal{E}(g^{\prime},g^{\prime})\geq 2\cdot\frac{|A|}{2^{n}}\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right)$
    for nonnegative $g$, where $g^{\prime}:=g/\mathbb{E}_{A}[g]$, and $\mathbb{E}_{A}[g]$ is the conditional expectation on $A$.
We generalize Samorodnitsky’s inequality to the $p$-biased setting so that the resulting functional inequality recovers the sharp constant in the $p$-biased edge-isoperimetric inequality. Let $f\in L^{2}(\{0,1\}^{n},\mu_{p}^{\otimes n})$, where $\mu_{p}$ is the Bernoulli measure on $\{0,1\}$ with $\mu_{p}(1)=p$ and $\mu_{p}(0)=1-p$. For $x\in\{0,1\}^{n}$, write $|x|=|\{i\in[n]:x_{i}=1\}|$, so that
Many applications (e.g., to the study of percolation [4], threshold phenomena in random graphs [26, 31], and hardness of approximation [11]) rely upon the use of the biased measure on the discrete cube. The definition of Dirichlet quadratic form of $g$ with respect to the biased measure is
Definition 1.2.
$x_{i}\leq y_{i},\ \forall i\in[n]\Rightarrow g(x)\leq g(y).$
The following ‘biased’ generalization of (2) for increasing subsets is considered folklore (see [15, 25]).
Theorem 1.3.
Let $A$ be an increasing subset of $\{0,1\}^{n}$ and let $0<p<1$. Then
equality holds in (4) if and only if $A$ is an increasing subcube.
Inspired by (1), it is natural to seek a functional inequality on the $p$-biased cube for general real-valued (increasing) functions that, in the special case of indicator functions, reduces to the $p$-biased edge-isoperimetric inequality with the sharp constant. We establishes such an inequality, which recovers (4) by substituting $g=\mathbbm{1}_{A}$.
Theorem 1.4.
Let $A$ be a subset of $\{0,1\}^{n}$ and let $g$ be a real-valued increasing function on $\{0,1\}^{n}$ supported on $A$. Then for any $0<p<1$,
Remark.
Note that the logarithmic Sobolev inequality (3) can be generalized to $p$-biased measure as follows, see [7, 9]:
$\mathcal{E}_{p}(f,f)\geq\frac{1}{p(1-p)}\cdot\frac{1-2p}{\ln(1-p)-\ln(p)}\cdot{\rm Ent}_{p}(f^{2}).$ (6)
When $f:=\mathbbm{1}_{A}$, we have the following
$p\cdot\mathcal{E}_{p}(f,f)\geq\frac{\ln(1/p)}{1-p}\cdot\frac{1-2p}{\ln(1-p)-\ln(p)}\cdot\mu_{p}(f)\log_{p}\mu_{p}(f),$ (7)
and hence for $p\to 0$,
$(1+o_{p}(1))p\cdot\mathcal{E}_{p}(f,f)\geq\mu_{p}(f)\log_{p}\mu_{p}(f),$
which again can not cover the original $p$-biased version of edge-isoperimetric inequality (4).
We view the $p$-biased hypercube as a weighted undirected graph $Q^{n}_{p}=(\{0,1\}^{n},E,w_{p})$, where for every vertex $x\in\{0,1\}^{n}$ and coordinate index $i\in[n]$, the edge weight is defined by
$\begin{split}w_{p}(\{x,x\oplus e_{i}\})&=p^{|x|}(1-p)^{n-|x|}+p^{|x\oplus e_{i}|}(1-p)^{n-|x\oplus e_{i}|}\\ &=p^{\sum_{j\in[n]:j\neq i}x_{j}}(1-p)^{n-\sum_{j\in[n]:j\neq i}x_{j}}\left(p^{x_{i}}(1-p)^{-x_{i}}+p^{x_{i}\oplus e_{i}}(1-p)^{-x_{i}\oplus e_{i}}\right)\\ &=p^{\sum_{j\in[n]:j\neq i}x_{j}}(1-p)^{n-\sum_{j\in[n]:j\neq i}x_{j}-1}.\end{split}$
Thus each edge is weighted proportionally to the combined $p$-biased mass of its two endpoints.
A random walk on a graph is a process that begins at some vertex, and at each time step moves to another vertex. When the graph is unweighted, the vertex the (simple) random walk moves to is chosen uniformly at random among the neighbors of the present vertex. When the graph is weighted, it moves to a neighbor with probability proportional to the weight of the corresponding edge. In [30], Samorodnitsky observed that the functional inequality (1) is equivalent to a probabilistic statement about random walks on the hypercube: subcubes maximize mean first exit time among all subsets of the same cardinality. Using the same argument, our new $p$-biased edge-isoperimetric inequality (5) admits an analogous probabilistic interpretation in terms of a simply random walk on the weighted hypercube $Q^{n}_{p}$.
Let $A\subseteq\{0,1\}^{n}$ and define the following process. Choose a uniformly random point $x\in A$ and then moving to one of its $n$ neighbors with probability ${\rm P}_{p}(x,x\oplus e_{i})=\omega_{p}(\{x,x\oplus e_{i}\})/d_{p}(x)$, where $d_{p}(x)=\sum_{i=1}^{n}\omega_{p}(\{x,x\oplus e_{i}\})$ is the weighted degree of vertex $x$. Let $Y$ be the time that it takes the walk to exit $A$ for the first time. We refer to $\mathbb{E}_{p}[Y]$ as the mean first exit time of $A$ under the $p$-biased random walk.
Corollary 1.5.
Increasing subcubes maximize mean first exit time among all increasing subsets of the weighted hypercube $Q_{p}^{n}$ of the same cardinality. More precisely, for any increasing subset $A\subset\{0,1\}^{n}$,
$\mathbb{E}_{p}[Y]\leq\frac{n}{\log_{p}(\mu_{p}(A))},$ (8)
equality holds if and only if $g$ is indicator functions of increasing subcubes.
1.2 Poincaré inequality on increasing sets
@@
Discrete-time random walks on finite graphs $G=(V,E)$ model stochastic processes where a walker iteratively transitions between adjacent vertices. Formally, at each timestep $t$, the walker moves from current vertex $u$ to neighbor $v$ with probability $P(u,v)=\frac{1}{{\deg}(u)}$, defining a Markov chain with state space $V$ and transition matrix $P$. The stationary distribution $\pi$, satisfying $\pi P=\pi$, is given by $\pi(v)=\frac{\deg(v)}{2|E|}$ for connected non-bipartite graph. The mixing time $t_{{\rm mix}}(\varepsilon)$ quantifies the convergence rate to stationarity, defined as the minimal $t$ such that:
@@
Ding and Mossel [10] introduced the study of the random walk on $\{0,1\}^{n}$ censored to $A$. This walk, when at state $x\in A$, attempts to flip a uniformly random coordinate; it moves to the neighbor if it resides in $A$, otherwise it stays at $x$. Note that if we do not impose any restrictions on $A$, the censored random walk may mix well, and also may mix extremely poorly even when $A$ is large and connected, see [17, Example 1,2]. A significant open question ([10, Question 1.1]) asked if the optimal mixing time belongs to $O_{c}(n\log n)$ holds for increasing subsets $A$ with constant density under the uniform measure.
Question 1.6 ([10, Question 1.1]).
@@
When the density $\mu(A)$ is constant, Ding and Mossel established a mixing time bound of $O(n^{3})$. Very recently, Fei and Ferreira Pinto Jr [17] show an improved mixing time bound of $O(n^{2})$ for constant-density increasing sets via establishing the following Poincaré inequality on the non-product space.
Theorem 1.7 ([17, Theorem 1.6]).
@@
Note that when $A=Q^{n}$, the above result recovers the classical Poincaré inequality on the hypercube. Their proof relies heavily on the connections between the directed heat flow and the directed Poincaré inequality. In this paper, we provide a simple inductive proof of their Poincaré inequality, though yielding a slightly worse constant (theirs being 1 versus ours at 2). Specifically, we prove the following theorem, a weaker version of Theorem 1.7.
@@
Nevertheless, we still obtain an $O(n^{2})$ mixing time bound for increasing sets $A$ of constant density via standard Markov chain theory ([27, Theorem 12.4]).
Corollary 1.9.

### v3

# Functional Inequalities and Random Walks on Increasing Subsets of the Hypercube

@@ Abstract
Motivated by random walks on subsets of the hypercube, we prove two discrete functional inequalities on the hypercube. First, we give a short, elementary proof of the Poincaré inequality on increasing subsets of the cube recently established by Fei and Ferreira Pinto Jr [fei2025spectral], which yields an $O(n^{2})$ upper bound on the mixing time of censored random walks, improving upon previous bounds. Second, adapting Samorodnitsky’s induction method [Samorodnitsky17] to the $p$-biased setting, we establish a sharp $p$-biased edge-isoperimetric inequality for real-valued increasing functions, which recovers the classic biased edge-isoperimetric inequality for increasing sets and identifies increasing subcubes as the extremizers. This result also admits a probabilistic interpretation in terms of maximizing the mean first exit time of biased random walks.
Keywords: Random walks, mixing time, Poincaré inequality, edge-isoperimetry, $p$-biased hypercube, mean first exit time
MSC codes: 06E30, 05C81, 60C05, 60G50
@@ 1 Introduction
Functional inequalities on the discrete hypercube/Hamming graph, such as the Poincaré, log-Sobolev, and edge-isoperimetric inequalities, form a central part of modern discrete analysis. They connect Boolean function analysis, discrete isoperimetric problems, and the spectral theory of Markov chains, providing powerful tools for studying concentration of measure, threshold phenomena, and mixing times of random walks. In the classical uniform product setting on the hypercube $\{0,1\}^{n}$, these inequalities are well understood: sharp constants are known (e.g., the exact log-Sobolev constant [BT2006logsobolev]) and elegant proofs arise via tensorization [gross1975logarithmic], semigroup methods [Ivanisvili2024KKL, ivanisvili2020rademacher], or discrete Fourier analysis [Eldan2025Isoperimetric]. In this paper, we depart from the product setting–by restricting to structured subsets or by working under biased measures, and aim at deriving sharp functional inequalities in these spaces and applying them to characterize the convergence behaviors of relevant random walks.
Beyond the product setting, the study of censored random walks on increasing subsets of the hypercube is motivated by foundational challenges in Markov Chain Monte Carlo (MCMC) for constrained sampling [Jer1998, guruswami2016rapidly, MT2006, LP2017]. A canonical example is sampling uniformly from configurations of a statistical mechanics model (e.g., critical percolation [DM2014]) that satisfy a non-local property such as a left-to-right crossing. While the ambient space is a product measure and trivial to sample from, the constraint set forms a highly correlated, nonproduct subset. Rejection sampling becomes statistically inefficient for rare events, prompting the use of local chains that remain within the constraint set. The natural censored dynamics re-randomizes one coordinate at a time and accepts the move only if the property is preserved. This process evolves on an increasing subset, and monotonicity ensures irreducibility. The central algorithmic question is to quantify the mixing time and the related survival/exit characteristics of this censored walk, since these quantities determine the efficiency and robustness of constrained sampling in combinatorics, statistical physics, and property testing.
1.1 Poincaré inequality and random walks on increasing sets
An illustration of non-product structure is motivated from censored random walks, where product structure is explicitly broken. If we censor the simple random walk on $\{0,1\}^{n}$ to a subset $A$ (only moves that stay in $A$ are accepted), the chain now lives on a non-product space with geometry dictated by the boundary of $A$. Currently, there is a lack of good estimation for the mixing time of such random walks. Let us introduce this problem in details.
@@ 1.1 Poincaré inequality and random walks on increasing sets
We work on the $n$-dimensional hypercube. The Hamming graph is the graph $Q^{n}=(\{0,1\}^{n},E)$ with $2^{n}$ vertices indexed by binary strings, where two vertices are adjacent (written $x\sim y$) iff they differ in exactly one coordinate. Discrete-time random walks on the Hamming graph (or more generally, on finite graphs $G=(V,E)$) model stochastic processes where a walker iteratively transitions between adjacent vertices. Formally, at each timestep $t$, the walker moves from current vertex $u$ to neighbor $v$ with probability $P(u,v)=\frac{1}{{\deg}(u)}$, defining a Markov chain with state space $V$ and transition matrix $P$. The stationary distribution $\pi$, satisfying $\pi P=\pi$, is given by $\pi(v)=\frac{\deg(v)}{2|E|}$ for connected non-bipartite graph. The mixing time $t_{{\rm mix}}(\varepsilon)$ quantifies the convergence rate to stationarity, defined as the minimal $t$ such that:
@@
Ding and Mossel [DM2014] introduced the study of the random walk on $\{0,1\}^{n}$ censored to $A$. This walk, when at state $x\in A$, attempts to flip a uniformly random coordinate; it moves to the neighbor if it resides in $A$, otherwise it stays at $x$. Note that if we do not impose any restrictions on $A$, the censored random walk may mix well, and also may mix extremely poorly even when $A$ is large and connected, see [fei2025spectral, Example 1,2]. A significant open question ([DM2014, Question 1.1]) asked if the optimal mixing time belongs to $O_{c}(n\log n)$ holds for increasing subsets $A$ with constant density under the uniform measure.
Question 1 ([DM2014, Question 1.1]).
@@
When the density $\mu(A)$ is constant, Ding and Mossel established a mixing time bound of $O(n^{3})$. Very recently, Fei and Ferreira Pinto Jr [fei2025spectral] make a breakthrough on this topic by showing an improved mixing time bound of $O(n^{2})$ for constant-density increasing sets. Their result is obtained via establishing the following Poincaré inequality on the non-product space.
Theorem 1.1 ([fei2025spectral, Theorem 1.6]).
@@
Note that when $A=Q^{n}$, the above result recovers the classical Poincaré inequality on the hypercube. Their proof relies heavily on the connections between the directed heat flow and the directed Poincaré inequality. In this paper, we provide a simple inductive proof of their Poincaré inequality, though yielding a slightly worse constant (theirs being 1 versus ours at 2). Specifically, we use the induction method to prove the following theorem, a weaker version of Theorem 1.1.
@@
Nevertheless, we still obtain an $O(n^{2})$ mixing time bound for increasing sets $A$ of constant density via standard Markov chain theory ([LP2017, Theorem 12.4]).
Theorem 1.3 ([fei2025spectral]).
@@
Fei and Ferreira Pinto Jr. [fei2025spectral] develop an $L^{2}$ functional–analytic framework for increasing sets $A\subseteq\{0,1\}^{n}$. They introduce a directed gradient that counts only order-consistent edges (down moves), which defines a directed Laplacian $L_{A}^{\to}$ and energy $\mathcal{E}_{A}^{\to}(f)$. They then run the monotone heat flow generated by $L_{A}^{\to}$ and prove an energy–variance dissipation inequality along the flow, yielding a directed Poincaré inequality with constant $C(\mu(A))=O(1)$ whenever $\mu(A)$ is bounded away from 0. A second ingredient is an approximate FKG inequality on increasing sets that supplies quantitative positive-correlation bounds. Together these tools give an optimal (up to constants) spectral gap for the censored walk on $A$.
In contrast, our method is designed to be simple and easy to check. We establish a slightly weaker, yet asymptotically optimal, Poincaré inequality by a classical induction-on-dimension argument. The proof restricts functions to lower-dimensional subcubes and uses the inherent increasing structure $A_{0}\subseteq A_{1}$ to decompose the variance and the Dirichlet form fiberwise, ultimately reducing the task to an explicit five-point inequality that can be verified directly. The resulting constant is within a factor $2$ of the optimal value and already implies the same mixing bound $O(n^{2})$ for censored walks. Because every step is combinatorial and closed-form, the proof is short, self-contained, and readily adaptable.
1.2 Biased Samorodnitsky’s inequality and exit time of random walks from increasing sets
Another illustration of non-product structure is motivated from characterizing the mean first exit time of a random walk from an increasing set.
Let $g:\{0,1\}^{n}\to\mathbb{R}$, we define the Dirichlet form of $g$ as
where the expectation is taken with respect to the uniform probability measure $\mu:=\mu_{1/2}^{\otimes n}$ on the hypercube. Furthermore, we denote by $x\sim A$ that $x$ is drawn uniformly at random from the subset $A\subseteq\{0,1\}^{n}$. In 2017, Samorodnitsky [Samorodnitsky17] established the following discrete functional inequality on the hypercube.
Theorem 1.4 (Samorodnitsky [Samorodnitsky17]).
When $g$ is an indicator function of $A$, Samorodnitsky’s inequality Eq. 5 implies the classic edge-isoperimetric inequality on $Q^{n}$: for every $A\subset\{0,1\}^{n}$,
$|\nabla A|\geq|A|\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right),$ (6)
where the edge boundary $\nabla(A)$ is the set of edges between $A$ and its complement $A^{c}=V(Q^{n})\setminus A$. Equality in Eq. 6 holds if and only if $A$ is a subcube of co-dimension $k$, for any $0\leq k\leq n$.
Besides the motivation from generalization of Haper’s edge-isoperimetric inequality, Samorodnitsky’s inequality admits a natural interpretation in languages of random walk on the hypercube. Specifically, it is equivalent to a sharp estimate of mean first exit time of a random walk from a subset with a given size [Samorodnitsky17].
We now generalize Samorodnitsky’s inequality to the $p$-biased setting so that the resulting functional inequality recovers the sharp constant in the $p$-biased edge-isoperimetric inequality. Let $f\in L^{2}(\{0,1\}^{n},\mu_{p}^{\otimes n})$, where $\mu_{p}$ is the Bernoulli measure on $\{0,1\}$ with $\mu_{p}(1)=p$ and $\mu_{p}(0)=1-p$. For $x\in\{0,1\}^{n}$, write $|x|=|\{i\in[n]:x_{i}=1\}|$, so that
Many applications (e.g., to the study of percolation [BKS1999], threshold phenomena in random graphs [KLMM2024globalhyper2024, Talagrand1994randomgraph], and hardness of approximation [DS2005Annal]) rely upon the use of the biased measure on the discrete cube. The definition of Dirichlet quadratic form of $g$ with respect to the biased measure is
Definition 1.
$x_{i}\leq y_{i},\ \forall i\in[n]\Rightarrow g(x)\leq g(y).$
The following ‘biased’ generalization of Harper’s edge-isoperimetric inequality for increasing subsets is considered folklore (see [EKL2019biased, Kahnkalai07]).
Theorem 1.5.
Let $A$ be an increasing subset of $\{0,1\}^{n}$ and let $0<p<1$. Then
equality holds in Eq. 7 if and only if $A$ is an increasing subcube.
Inspired by Eq. 5, it is natural to seek a functional inequality on the $p$-biased cube for general real-valued (increasing) functions that, in the special case of indicator functions, reduces to the $p$-biased edge-isoperimetric inequality with the sharp constant. We establishes such an inequality, which recovers Eq. 7 by substituting $g=\mathbbm{1}_{A}$.
Theorem 1.6.
Let $A$ be a subset of $\{0,1\}^{n}$ and let $g$ be a real-valued increasing function on $\{0,1\}^{n}$ with support $A$. Then for any $0<p<1$,
We view the $p$-biased hypercube as a weighted undirected graph $Q^{n}_{p}$, where for every vertex $x\in\{0,1\}^{n}$ and coordinate index $i\in[n]$, the edge weight is defined by
$\begin{split}w_{p}(\{x,x\oplus e_{i}\})&=\mu_{p}^{\otimes n}(x)+\mu_{p}^{\otimes n}(x\oplus e_{i})=\mu_{p}^{\otimes(n-1)}(x_{\backslash i}),\end{split}$
where $x_{\backslash i}=(x_{1},...,x_{i-1},x_{i+1},...,x_{n})$. Thus each edge is weighted proportionally to the combined $p$-biased mass of its two endpoints.
A random walk on a graph is a process that begins at some vertex, and at each time step moves to another vertex. When the graph is unweighted (i.e., $p=1/2$), the vertex the (simple) random walk moves to is chosen uniformly at random among the neighbors of the present vertex. When the graph is weighted, it moves to a neighbor with probability proportional to the weight of the corresponding edge. In [Samorodnitsky17], Samorodnitsky observed that the functional inequality Eq. 5 is equivalent to a probabilistic statement about random walks on the hypercube: subcubes maximize mean first exit time among all subsets of the same cardinality. Using the same argument, our new $p$-biased edge-isoperimetric inequality Eq. 8 admits an analogous probabilistic interpretation in terms of a simply random walk on the weighted hypercube $Q^{n}_{p}$.
Let $A\subseteq\{0,1\}^{n}$ and define the following random walk $X_{t}$. The random walk starts at a random point $X_{0}=x\in A$ following from the $p$-biased conditional law, i.e., $\nu(x):=\frac{\mu_{p}(x)}{\mu_{p}(A)},\forall x\in A$, and at each step, moves to one of its $n$ neighbors with probability ${\rm P}_{p}(x,x\oplus e_{i})=\omega_{p}(\{x,x\oplus e_{i}\})/d_{p}(x)$, where $d_{p}(x)=\sum_{i=1}^{n}\omega_{p}(\{x,x\oplus e_{i}\})$ is the weighted degree of vertex $x$. Let $Y$ be the (random) time that it takes the walk to exit $A$ for the first time for the $p$-biased walk. We refer to $T(A,p)=\mathbb{E}[Y]$ as the mean first exit time of $A$ under the $p$-biased walk. We further introduce the weighted mean first exit time as
$\tilde{T}(A,p):=\mathbb{E}\left[\sum\limits_{t=0}^{Y-1}\phi(X_{t})\right],{\rm where}\ \phi(x):=\frac{\mu_{p}(x)}{d_{p}(x)}.$
Theorem 1.7.
Let $A\subset\{0,1\}^{n}$ be increasing and $0<p<1$. Then
  • (1)
    $\tilde{T}(A,p)\leq\frac{p}{\log_{p}\mu_{p}(A)}$ with equality if and only if $A$ is an increasing subcube.
  • (2)
    $T(A,p)\leq\frac{p}{\min\{p,1-p\}}\cdot\frac{n}{\log_{p}(\mu_{p}(A))}$ with equality if and only if $p=\tfrac{1}{2}$ and $A$ is an increasing subcube.
Example 1 (Mean first exit time for the $p$-biased walk on increasing subcubes).
Fix $0<p<1$ and $n\in\mathbb{N}$. Let $C:=\{x\in\{0,1\}^{n}:x_{1}=\cdots=x_{k}=1\}$ be the increasing subcube of codimension $k$, i.e., $k=\log_{p}\mu_{p}(A)$. Then
$T(C,p)=1+2p\left(\frac{n}{k}-1\right).$
In particular, for $p=\tfrac{1}{2}$ this reduces to $T(C,p)=n/k$.
1.3 Organization